What is the solubility of Cr(OH)₃ at a pH of 11.10? (Ksp Cr(OH)₃ is 6.70 × 10⁻³¹)
Answer #1

Crouz is abase pH + pon = 14 pon- 14-pH - = 14 - 11.10 = 2.90 [ou] = 10-pon pole-log[con 1) >> Con ] - 10-2.90 - 1.26310-3 S